solve the system of equation by graphing
x+y=8
x-y=8

Answer by mananth(16946)

1 x + 1 y = 8 .............1
-1 x -1 y = 8 .............2

equation I
when x= 0 y= 8 ,( 0 , 8 )
when y = 0 x= 8 ,( 8 , 0 )
Equation 2
when x= 0 y= -8 ( 0 , -8 )
when y = 0 x= 8 ( 0 , 8 )


The solution is (8,0)